OpenStudy (anonymous):

can some tell me how to solve this problem? -36c^12 over 12c^6

OpenStudy (amistre64):

top: -36 c^12 bottom: 12 c^6 (-36/12) reduces to -3 c^(12-6) = c^6 Answer: -3c^6

OpenStudy (amistre64):

when multiplying exponents: a^3 * a^5; you add the exponents together: a^(3+5) = a^8 When you have an exponent on the bottom of a fraction: 1/(a^3) You can re-write it as: 1/(a^3) = a^(-3) Lets add try this: (a^4)/(a^3) = (a^4) * (a^-3) = a^(4-3) = a^1

OpenStudy (anonymous):

-36c^12/12c^6 =-3c^12-6=-3c^6